Assume that, on January 1, 2016, Matsui Co. paid $1,296,000 for its investment in 48,000 shares of Yankee Inc. Further, assume that Yankee has 240,000 total shares of stock issued. The book value and fair value of Yankee's identifiable net assets were both $480,000 at January 1, 2016. The following information pertains to Yankee during 2016:

  

  Net income $240,000
  Dividends declared and paid $72,000
  Market price of common stock on 12/31/2016 $29/share

  

What amount would Matsui report in its year-end 2016 balance sheet for its investment in Yankee?

  

$1,608,000.

$1,368,000.

$1,329,600.

None of these answer choices is correct.
Answer #1

$$ \begin{aligned} \text { Percentage of Holding } &=\frac{48,000}{240,000} \times 100 \\ &=20 \% \end{aligned} $$

Value of Investment \(=\) Cost of investment \(+\) Share in Net income - Share in dividend

$$ \begin{array}{l} =\$ 1,296,000+(\$ 240,000 \times 20 \%)-(\$ 72,000 \times 20 \%) \\ =\$ 1,296,000+\$ 48,000-\$ 14,400 \\ =\$ 1,329,600 \end{array} $$

Hence, the correct option is \(\mathbf{\$ 1 , 3 2 9 , 6 0 0}\).